OpenStudy (anonymous):

Which is equal to 7 2/3 7 1/6 7 1/3 7 3/2

OpenStudy (anonymous):

missing something?

OpenStudy (anonymous):



OpenStudy (anonymous):

ooooh ok

OpenStudy (anonymous):

and the answers are with exponents right? not for example mixed numbers

OpenStudy (anonymous):

in a fractional exponent the numerator is the power and the denominator is the root i.e \[\huge \sqrt[n]{x^m}=x^\frac{m}{n}\]

OpenStudy (anonymous):

in your case \(n=2,m=3\) and of course \(x=7\)

OpenStudy (anonymous):

so it would be the first one

OpenStudy (anonymous):

no you have it backwards

OpenStudy (anonymous):

oh the last one

OpenStudy (anonymous):

the root goes in the denominator, the power in the exponent yes, the last one